Lepton + Jets Reconstruction

In this mode, have measured almost all t t daughters

• Lepton 4-momentum solid and νl is missing only pz

• W → qq̄′ observed as 2 jets

• bottom quarks observed as jets (perhaps tagged)

Problems:

• QCD radiation generates additional jets

• Combinatorics in identifying jets ↔ partons

• Background events

Jet-Parton Corrections

Have to first correct observed E_{T} to parton E_{T}

• Corrections determined by MC and testbeam studies • Make additional corrections for t t topology

Tests of CDF Reconstruction

Can test how well procedure works with MC events

• Look at correct and “best” assignments

• Also look at how background would reconstruct

Background Effects

Primary backgrounds are from W bb̄ and fake leptons

• Use background rates from counting experiments

• Can check W bb̄ shape by looking at Z+ jets

• Use “non-isolated” lepton + jet sample for fakes

• Look identical to VECBOS W + jets prediction ⇒ Use VECBOS W + jets MC for bkg shape

CDF Mass Results

Estimate background based on W +jet sample

• Use known fake tag rates and W bb̄

• Background shape in Mtop comes from MC calculation

• Estimate background to be 6.9{+3.4 \atop −2.4} events – Primarily QCD W+jet

Mtop = 176 ± 8 ± 10 GeV/c2

Evidence for W → qq̄

Observation of W → qq̄ is important

• Strong evidence of signal in W + W −bb̄ channel • Can calibrate jet energy scale

Have to refit data relaxing W mass constraint

• Look for W signal plotting all mass combinations – Makes it tough since 6 combos/event

• Reduce backgrounds by requiring H > 300 GeV

CDF BR Results

Result from dilepton data is R = 0.875 ± 0.4

Result from lepton + jets data is R = 0.87 ± 0.3

t t Invariant Mass

Data is statistics limited

• There is no apparent enhancement

• An upper limit on Top-Color Z′ mass and cross section coming soon